MOA-II Obsevation Log Monthly Summary 2007/10
Return to index
Previous month
Next month
TARGET | RUN | TOTAL RUN | TOTAL FIELD | TOTAL EXPOSURE (sec) |
GB | 11141-11382 | 242 | 944 | 56640 |
LMC | 1802-1896 | 95 | 474 | 135300 |
SMC | 445-456 | 12 | 24 | 7080 |
ALERT | - | 0 | 0 | 0 |
GRB | - | 0 | 0 | 0 |
TOTAL | --- | 349 | 1442 | 199020 (55h17m00s) |
GB FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
37 | 42 | 34 | 29 | 123 | 25 | 37 | 29 | 133 | 136 | 52 | 33 |
Field13 | Field14 | Field15 | Field16 | Field17 | Field18 | Field19 | Field20 | Field21 | Field22 | Field23 |
27 | 25 | 22 | 28 | 24 | 22 | 22 | 19 | 18 | 27 | 0 |
LMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
17 | 15 | 15 | 14 | 13 | 19 | 90 | 88 | 82 | 28 | 20 | 20 |
Field13 | Field14 | Field15 | Field16 |
26 | 27 | 0 | 0 |
SMC FIELD STATISTICS
Field1 | Field2 | Field3 | Field4 | Field5 | Field6 | Field7 | Field8 | Field9 | Field10 | Field11 | Field12 |
12 | 12 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 |
SPECIAL TARGET NAME
CCD TEMPERATURE
Previous month
Next month

VACUUM
Previous month
Next month

TELESCOPE TEMPERATURE
Previous month
Next month

COMMENTS
2007/10/1: D23739 (60s) / night time = 34140(sec)
2007/10/2: / night time = 33900(sec)
2007/10/3: D2743 (300s) / night time = 33720(sec)
2007/10/4: D2755(300sec) is OK, wind meter in B&C was broken / night time = 33480(sec)
2007/10/5: No observation / night time = 33240(sec)
2007/10/6: D2756(60 sec) is OK / night time = 33000(sec)
2007/10/7: D2758(60sec) ,D2759(300sec) are OK ,bad seeing / night time = 32820(sec)
2007/10/8: D2761(60sec) D2763(300sec) are OK / night time = 32580(sec)
2007/10/9: D2768(60sec) ,D2775(300sec) are OK / night time = 32400(sec)
2007/10/10: No observation / night time = 32160(sec)
2007/10/11: No observation,Power failure (9:30-10:30am) ,Pumping on going / night time = 31980(sec)
2007/10/12: Took gb10 once every two exposure for OGLE514(High mag event),D2777(60sec),D2778(300sec) are OK,Pump off before observation / night time = 31740(sec)
2007/10/13: D2779(60sec),D2780(300sec) are OK ,Took gb10 intensively for OGLE514(MOA464) / night time = 31440(sec)
2007/10/14: Took gb10 once every two exposure for MOA464(OGLE514),D2786(60sec),D2814(300sec) chip2 noisy / night time = 31260(sec)
2007/10/15: Took gb10 as often as gb5,9 for MOA464(OGLE514) , D2820(60sec),D2821(300sec) are OK / night time = 31020(sec)
2007/10/16: No observation,FLAT(F1989-F2024),D2822(15sec) / night time = 30780(sec)
2007/10/17: No observation / night time = 30600(sec)
2007/10/18: gb22 was skipped due to the MOON,DARK 2830(60sec),2831(300sec) / night time = 30360(sec)
2007/10/19: DARK 2839(300sec) / night time = 30120(sec)
2007/10/20: No observation / night time = 29880(sec)
2007/10/21: DARK 2842(300sec) / night time = 29640(sec)
2007/10/22: No observation,Pumping on going / night time = 29400(sec)
2007/10/23: Pump OFF before obserbation,DARK 2843(60sec),2850(300sec) / night time = 29160(sec)
2007/10/24: The Moon was bright,DARK 2873(60sec),2874(200sec),2875(240sec),2876(300sec) / night time = 28920(sec)
2007/10/25: Vacuum sensor broken,No observation / night time = 28680(sec)
2007/10/26: No observation,Change vacuum sensors ,Pumping on going / night time = 28440(sec)
2007/10/27: Exposure time for LMC and SMC is 240sec due to the Moon,DARK 2878(60sec),2880(240sec) / night time = 28260(sec)
2007/10/28: DARK 2883(60sec), 2884(300sec) are OK. Pumping on going. / night time = 28020(sec)
2007/10/29: Change exposure time 240sec,DARK 2885(240sec),2891(300sec), Pumping off / night time = 27780(sec)
2007/10/30: DARK D2895(300sec) is OK. / night time = 27480(sec)
2007/10/31: DARK D2900(300sec) is OK. Pumping on going. / night time = 27240(sec)
TOTAL NIGHT TIME: 953640(sec)
OPERATING RATE: 26.9 %
GB: 9.9 %
LMC: 16.2 %
SMC: 0.8 %
ALERT: 0.0 %
GRB: 0.0 %
Return to index
Previous month
Next month